OpenStudy (anonymous):

Solve the system shown. Put your answer in (x, y) format with no spaces. y + 4x = 8 5x + 2y = 13

OpenStudy (anonymous):

Are both problem's variables supposed to have the same value? Is that was you've been taught in class to do?

OpenStudy (anonymous):

Next time, please put these problems in separate questions OR number them. 1. y + 4x = 8 8/4 = 2 2/1 = 2 (1 is actually y but is represented as 1 to show that it is 1y) 2. 5x + 2y = 13 13 - 5 = 8 8/2 = 4

OpenStudy (anonymous):

Correct? Idk.........

OpenStudy (anonymous):

wait then shouldn't the Asker have included that?